OpenStudy (anonymous):

Through:(1,5),slope=4

OpenStudy (anonymous):

I assume you want an equation with that slope that goes through that point. Recall the point slope formula y - y1 = m(x - x1) (x1, y1) is any point on the line m is your slope

OpenStudy (nirmalnema):

use one point form. (y-y1)=slope (x-x1)

OpenStudy (anonymous):

Use this : Y2-Y=m(x2-x) m is slope = 4 Y2=5 x2=1 and x and y are in equation :) Got it ?:)

OpenStudy (anonymous):

ok...we've given enough help. if they don't respond...stop helping.

OpenStudy (nirmalnema):

y-5=4(x-1) y-5=4x-4 4x-y-1